OpenStudy (anonymous):

Find the derivatives of f(g(x)) and g(f(x)) f(u)=u^3, g(x)= 1/1+x

OpenStudy (anonymous):

f= 3u^2 g= -1/(1-X)^2 i would think

OpenStudy (mertsj):

Satellite appears and the crowd gathers.

OpenStudy (anonymous):

just spacing out

OpenStudy (anonymous):

So you just find the derivatives and you don't have to use the chain rule?

OpenStudy (anonymous):

nvm idk dont you mean x instead of u for the f(u)= U^3

OpenStudy (anonymous):

\[f(g(x))=f(\frac{1}{1+x})=(\frac{1}{1+x})^3\] \[f(g(x))'=3(\frac{1}{1+x})^2\times \frac{-1}{(1+x)^2}\]

OpenStudy (anonymous):

yes by the chain rule

OpenStudy (anonymous):

yeah satellite is correct i was simply using wut u game me w/o subbing in for the f(g(x)) part

OpenStudy (anonymous):

\[g(f(x))=g(x^3)=\frac{1}{1+x^3}\] \[g(f(x))'=-\frac{1}{(1+x^3)^2}\times 3x\]

OpenStudy (raspberryjam):

Essentially you are finding the derivative of \[f(g(x))=(\frac{1}{1+x})^3\] and \[g(f(x))=(\frac{1}{1+x^3})\]

OpenStudy (anonymous):

both clean up a tiny bit with some algebra

OpenStudy (anonymous):

Ah makes sense. I see where I went wrong. Thanks everyone.

OpenStudy (anonymous):

hey satellite are u good with calc 2? i got a few questions